OpenStudy (anonymous):

y varies inversely with x, and x = 88 when y = 11. Find the inverse variation equation and use it to find the value of y when x = 4.

OpenStudy (zzr0ck3r):

y = k/x for some k solve for k with the given point 11 = k/88 k = 11*88 = 968 so y = 968/x let x = 4 y = 968/4

OpenStudy (anonymous):

its 242

OpenStudy (anonymous):

i think thanks